A Complete Bibliography of Publications in Markup Languages: Theory & Practice

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

07 August 2013 Version 1.06

Title word cross-reference

RM [32].

2nd [106].

7 [91].

acquired [55]. al. [92, 98]. analysis [11]. analytical [81]. Annotated [7, 14, 24, 35]. Anomalous [27]. Applications [80]. approach [84, 91]. architectural [77]. Architecture [23, 54]. Architectures [39, 94]. Author [9, 16, 26, 37]. Authoring [42, 84]. Ayres [92].

based [19]. Best [38]. Best-Practice [38]. better [60]. between [71]. Beyond [56, 82].

Biographies [9, 16, 26, 37]. Black [24]. BNC [11]. Bob [24, 100]. Book [24, 33, 34, 44]. Books [8, 15, 25, 36]. border [54]. Bradley [24, 96]. Budget [21]. Building [60, 80]. Burke [97].

Cagle [98]. Calculus [78]. Can [3, 45]. case [11]. Caterpillars [43]. checking [46]. Cheryl [24]. classifications [55]. client [73]. code [49]. collections [13]. Commentary [17]. Companion [24, 96]. Complexity [79]. computer [49]. Conformance [30, 75]. consistently [45]. Consultant [67]. Content [3, 12, 42]. Contents [7, 14, 35]. Context [43, 79, 85]. Context-Free [79]. contract [77]. control [63]. Cookbook [14]. Corporate [29]. Crane [35, 99]. Creating [19]. critical [71]. criticizing [71]. Current [69].

Danny [92]. Data [27, 39, 48, 92, 104]. Dates [28, 46]. death [73]. define [57]. Definitive [59]. Demonstrational [47]. descriptive [65, 82]. descriptive/ procedural [65]. Developing [7, 38]. digital [51]. Discovering [31]. distinction [65, 82]. do [32]. DocBook [59]. Document [2, 39, 53, 57, 91]. Documents [5–7, 51]. Does [5]. DOM [83]. Doug [109]. Down [40]. draft [30]. draft-19990728-1650 [30]. DTD [70]. DTDs [7, 66, 71]. DuCharme [24, 100]. Dummies} [14].

Eddy [24]. Edge [108]. Edition [3, 62, 106, 108]. editions [71]. editors [73]. electronic [10]. engine [81, 83]. engineering [81]. English [24]. Entity [23]. Eric [97]. Essentials [101]. European [21]. Example [33]. Exceptions [79]. Expectation [10]. Experiences [45]. Expression [28]. Expressions [22, 46, 84].

faces [68]. family [66]. Ferrets [81]. Fitzgerald [101]. flawed [65]. flow [88]. Flynn [34]. formal [50]. Formalism [78]. forms [77]. Free [79]. FRESS [2]. functionalities [64]. Fung [102].

Gardner [103]. general [71]. Generation [4, 47, 72–74]. generator [87]. glasses [90]. Gnome [83]. Go [5]. Golem [60]. Graham [44]. Grammars [79]. Guide [44, 59, 103, 104]. Guidelines [38].

Handling [34]. Haystacks [31]. Helps [19]. Hierarchies [41]. Historical [76]. Hjelm [104]. Holzner [105]. HTML [102]. Hypertext [2, 64].

Ian [44]. IBM [19]. Imbalance [69]. Inadequacy [79]. Information} [14]. integrated [51]. Interactive [3]. interface [47]. interpretation [52]. interrelated [13]. Interview [17]. Introducing [31].

Java [58, 92, 97]. Jeni [108]. Joan [17]. John [103, 104, 107]. Journal [3]. Just [107].

Kay [106]. keep [66]. Khun [102].

Knowledge [29, 81]. Krik [24]. Kurt [98].

L [103]. Language [35, 87, 99]. Languages [1, 78]. Leonard [59]. Liam [44]. library [51]. linguistic [11]. literary [62]. Localization [93]. Logic [78, 91]. Ltd. [35, 99].

M [97]. maintain [66]. management [54]. Managing [13, 51, 53]. Mann [62]. many [68]. Map [31]. Maps [41, 55, 61, 63, 81, 89, 90]. Marked [6, 49]. Marked-Up [6, 49]. Markup [1, 2, 7, 18, 52, 69]. McGrath [33]. Meaning [52]. meets [10]. Memory [29]. messaging [54]. Michael [101, 106]. Model [7]. Models [12, 79]. modules [13]. Moultis [24]. MPEG [91]. MPEG-7 [91]. Muellner [59]. Muench [80]. Multimedia [78]. multiple [55].

Named [23]. Natanya [24]. Navigable [55]. Navigating [31]. Needles [31]. Neil [24, 96]. News [3]. Next [72–74]. next-generation [73]. Norman [59]. Note [13].

OASIS [75, 86]. one [70]. Onions [40]. Opinion [17]. Oracle [80]. Orlando [45]. overlaying [55].

Parsing [22]. Path [35, 78, 99]. patterns [50]. EDI [38]. procedural [65]. Xpath [75]. Perspective [76]. Peter [34, 104]. Pitts [24]. Pitts-Moultis [24]. Plain [24]. Planning [62]. Practical [34, 35, 99]. Practice [1, 13, 38]. Predicate [78, 91]. Primer} [14]. Problem [27]. procedural [82]. Processing [58, 88]. Professional [92, 98]. program [49]. Programmer [106]. Programming [6, 49]. Programs [34]. Project [10, 19, 21, 29, 45, 62]. property [77]. propositions [61]. publishing [10].

Quality [5]. queries [91]. Query [78]. Quickly [100]. Quin [44].

RDF [89, 90]. reality [10]. Received [8, 15, 25, 36]. Recipes [14]. Recognition [23]. Reference [106]. regarding [61]. Regular [22, 28, 46]. related [66]. relationship [71]. relationships [53]. Rendon [103]. Report [10, 19, 21, 29, 38]. Reports [30, 31]. Repurposing [3]. Reuse [3]. Review [44, 59, 80, 92]. Reviews [33, 34]. REX [22]. Robert [103].

S [44]. Sandra [24]. schemas [56]. Seán [33]. self [63]. self-control [63]. semantic [55, 89]. Semantics [50, 95]. semistructured [48]. set [77]. SGML [4, 5, 7, 10–12, 14, 21, 34, 72, 74, 76, 79]. Shallow [22]. simple [77]. Simpson [107]. Sites [18]. Smith [17]. society [10]. Softwrights [35, 99]. source [49]. specific [71]. Specification [24, 43, 44, 85]. Squib [20, 28, 40]. Standard [31]. Standards [30, 31, 38, 86]. Stark [104]. Steve [80]. Steven [105]. Street [3]. Structure [2, 49]. Structured [14, 18, 34, 39]. structures [53]. Structuring [7, 29]. Style [19]. stylesheet [47]. switchboard [66]. Syntactic [95]. Syntax [95]. System [2, 19, 21, 39]. systems [54].

Table [7, 14, 35]. tag [45]. Tale [18]. TaskGuide [19]. team [45]. technical [10]. Technique [43]. Techniques [13]. TEI [71]. Templates [41]. Tennison [108]. testing [75]. Text [7, 13, 34]. them [66]. Theory [1]. Thomas [62]. Tidwell [109]. Tool [42]. Toolkit [67]. Tools [4, 23, 34]. Topic [31, 41, 55, 61, 63, 81, 89, 90]. Topology [41].

Transcoding [42]. Transformation [35, 87, 88, 99]. Transformations [35, 99, 103]. Transforming [104]. Two [18, 61]. Type [41, 62]. types [57].

Ultimate [104]. UML [57].

Understanding [34]. Unicode [32]. Union [21]. Unlimited [108]. Updated [67]. use [32]. Using [11, 35, 49, 57, 58, 66, 99].

visual [84, 85]. vs [82].

Wall [3]. Walsh [59]. Web

[18, 42, 53, 89, 95, 104]. Welcome [1]. Where [5]. Whither [20]. Wizard [19]. Wizard-Style [19]. work [86]. Working [102]. World [94].

XML [7, 12, 14, 19, 22–24, 30, 33–

35, 38, 44, 48, 49, 51, 54, 57, 58, 64, 73, 78– 80, 87, 88, 90, 91, 93, 94, 99, 102, 103]. XML-based [19]. XML/EDI [38]. XPath [35, 50, 84, 99, 103, 108]. XSL [35, 96, 98, 99, 101, 107]. XSLT [35, 50, 68, 75, 87, 97, 99, 100, 102– 106, 108, 109]. XSLT/Xpath [75]. XSLTM [47]. XTL [87].

Yee [102].

Zarella [103].

References

Sperberg-McQueen:1999:WML

[1] C. M. Sperberg-McQueen and B. Tommie Usdin. Welcome to Markup Languages: Theory & Practice. Markup Languages: Theory & Practice, 1(1):1– 6, Winter 1999. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

                                             tronic). URL http://mitpress.mit.
                                             edu/journals/MLANG/us-sp.pdf.

DeRose:1999:DSM

[2] Steven J. DeRose and Andries van Dam. Document structure and markup in the FRESS hypertext system. Markup Languages: Theory & Practice, 1(1):7– 32, Winter 1999. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Karben:1999:NYC

[3] Alan Karben. News you can reuse: Content repurposing at The Wall Street Journal Interactive Edition. Markup Languages: Theory & Practice, 1(1):33– 45, Winter 1999. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

tronic). URL http://mitpress.mit.
edu/journals/MLANG/karben.pdf.

Matzen:1999:NGT

[4] R. W. Matzen. A new generation of tools for SGML. Markup Languages: Theory & Practice, 1(1):47– 74, Winter 1999. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

tronic). URL http://mitpress.mit.
edu/journals/MLANG/matzen.pdf.

Ramalho:1999:SDW

[5] José Carlos Ramalho, José João Almeida, Jorge Gustavo Rocha, and Pedro Henriques. SGML documents: Where does quality go? Markup Languages: Theory & Practice, 1(1):75– 90, Winter 1999. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Wood:1999:PMD

[6] Lauren Wood. Programming marked-up documents. Markup Languages: Theory & Practice, 1(1):91–100, Winter 1999.

CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Lapeyre:1999:ATCa

[7] Deborah A. Lapeyre. Annotated table of contents: Developing SGML DTDs From Text to Model to Markup; Structuring XML Documents. Markup Languages: Theory & Practice, 1(1):113– 118, Winter 1999. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Anonymous:1999:BRa

[8] Anonymous. Books received. Markup Languages: Theory & Practice, 1(1): 118, Winter 1999. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Anonymous:1999:ABa

[9] Anonymous. Author biographies. Markup Languages: Theory & Practice, 1(1):119–120, Winter 1999. CO-DEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Fahrenholz-Mann:1999:PRS

[10] Sally Fahrenholz-Mann. Project report: SGML for electronic publishing at a technical society: Expectation meets reality. Markup Languages: Theory & Practice, 1(2):1–30, Spring 1999. CO-DEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Burnard:1999:USL

[11] Lou Burnard. Using SGML for linguistic analysis: the case of the BNC. Markup Languages: Theory & Practice, 1(2):31– 51, Spring 1999. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Kilpelainen:1999:SXC

[12] Pekka Kilpeläinen. SGML & XML content models. Markup Languages: Theory & Practice, 1(2):53–76, Spring 1999. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Streich:1999:PNT

[13] Robert Streich. Practice note: Techniques for managing collections of interrelated text modules. Markup Languages: Theory & Practice, 1(2):77– 94, Spring 1999. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Anonymous:1999:ATC

[14] Anonymous. Annotated table of contents: XML for Dummies; The XML & SGML Cookbook: Recipes for Structured Information; XML: A Primer. Markup Languages: Theory & Practice, 1(2):95– 104, Spring 1999. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Anonymous:1999:BRb

[15] Anonymous. Books received. Markup Languages: Theory & Practice, 1(2): 105, Spring 1999. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Anonymous:1999:ABb

[16] Anonymous. Author biographies. Markup Languages: Theory & Practice, 1(2):106, Spring 1999. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Smith:1999:COI

[17] Joan M. Smith, C. M. Sperberg-McQueen, and B. Tommie Usdin. Commentary and opinion: Interview with Joan Smith. Markup Languages: Theory & Practice, 1(3):1–6, Summer 1999. CO-DEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Lubell:1999:SMW

[18] Joshua Lubell. Structured markup on the Web: a tale of two sites. Markup Languages: Theory & Practice, 1(3):7– 22, Summer 1999. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Tidwell:1999:PRI

[19] Doug Tidwell. Project report: IBM’s TaskGuide: An XML-based system for creating wizard-style helps. Markup Languages: Theory & Practice, 1(3):23– 39, Summer 1999. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Graham:1999:SW

[20] Tony Graham. Squib: Whither &? Markup Languages: Theory & Practice, 1(3):40, Summer 1999. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Catteau:1999:PRS

[21] Tom Catteau. Project report: An SGML system for the budget of the European Union. Markup Languages: Theory & Practice, 1(3):41–59, Summer 1999. CO-DEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Cameron:1999:RXS

[22] Robert D. Cameron. REX: XML shallow parsing with regular expressions. Markup Languages: Theory & Practice, 1(3):61–88, Summer 1999. CO-DEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Mikheev:1999:XTA

[23] Andrei Mikheev, Claire Grover, and Marc Moens. XML tools and architecture for named entity recognition. Markup Languages: Theory & Practice, 1(3):89–113, Summer 1999. CO-DEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Piez:1999:XCN

[24] Wendell Piez. The XML Companion, by Neil Bradley; XML: The Annotated Specification, by Bob DuCharme; XML In Plain English, by Sandra E. Eddy; The XML Black Book, by Natanya Pitts-Moultis and Cheryl Krik. Markup Languages: Theory & Practice, 1(3):114– 117, Summer 1999. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Anonymous:1999:BRc

[25] Anonymous. Books received. Markup Languages: Theory & Practice, 1(3): 118, Summer 1999. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Anonymous:1999:ABc

[26] Anonymous. Author biographies. Markup Languages: Theory & Practice, 1(3):119, Summer 1999. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Birnbaum:1999:PAD

[27] David J. Birnbaum and David A. Mundie. The problem of anomalous data. Markup Languages: Theory & Practice, 1(4):1–19, Fall 1999. CO-DEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Sperberg-McQueen:1999:SRE

[28] C. M. Sperberg-McQueen. Squib: Regular expression for dates. Markup Languages: Theory & Practice, 1(4):20– 26, Fall 1999. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Attipoe:1999:PRK

[29] Alfred Attipoe. Project report: Knowledge structuring for corporate memory. Markup Languages: Theory & Practice, 1(4):27–36, Fall 1999. CO-DEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Holman:1999:SRX

[30] G. Ken Holman. Standards reports: XML conformance (draft-19990728-1650). Markup Languages: Theory & Practice, 1(4):37–45, Fall 1999. CO-DEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Pepper:1999:SRN

[31] Steve Pepper. Standards reports: Navigating haystacks and discovering needles: Introducing the new topic map standard. Markup Languages: Theory & Practice, 1(4):47–74, Fall 1999. CO-DEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Graham:1999:UWI

[32] Tony Graham. Unicode^{RM}: What is it and how do I use it? Markup Languages: Theory & Practice, 1(4):75–102, Fall 1999. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Flynn:1999:BRX

[33] Peter Flynn. Book reviews: XML by Example, by Seán McGrath. Markup Languages: Theory & Practice, 1(4): 103–107, Fall 1999. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Mah:1999:BRU

[34] Carole E. Mah. Book reviews: Understanding SGML and XML Tools: Practical Programs for Handling Structured Text, by Peter Flynn. Markup Languages: Theory & Practice, 1(4):103– 107, Fall 1999. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Lapeyre:1999:ATCb

[35] Deborah A. Lapeyre. Annotated table of contents: Practical Transformation Using XSLT and Xpath (XSL Transformations and the XML Path Language), by Crane Softwrights Ltd. Markup Languages: Theory & Practice, 1(4):107– 110, Fall 1999. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Anonymous:1999:BRd

[36] Anonymous. Books received. Markup Languages: Theory & Practice, 1(4): 111, Fall 1999. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Anonymous:1999:ABd

[37] Anonymous. Author biographies. Markup Languages: Theory & Practice, 1(4):112, Fall 1999. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Bryan:2000:SRD

[38] Martin Bryan. Standards report: Developing best-practice guidelines for XML/ EDI. Markup Languages: Theory & Practice, 2(1):1–10, Winter 2000. CO-DEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Arnold-Moore:2000:SAS

[39] Timothy Arnold-Moore, Michael Fuller, and Ron Sacks-Davis. System architectures for structured document data. Markup Languages: Theory & Practice, 2(1):11–39, Winter 2000. CO-DEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Maler:2000:SO

[40] Eve Maler. Squib: Down with onions. Markup Languages: Theory & Practice, 2(1):40–44, Winter 2000. CO-DEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Rath:2000:TMT

[41] Hans Holger Rath. Topic maps: Templates, topology, and type hierarchies. Markup Languages: Theory & Practice, 2(1):45–64, Winter 2000. CO-DEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Hori:2000:ATW

[42] Masahiro Hori, Kouichi Ono, Goh Kondoh, and Sandeep Singhal. Authoring tool for Web content transcoding. Markup Languages: Theory & Practice, 2(1):65–80, Winter 2000. CO-DEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Bruggemann-Klein:2000:CCS

[43] Anne Bruggemann-Klein and Derick Wood. Caterpillars: a context specification technique. Markup Languages: Theory & Practice, 2(1):81–106, Winter 2000. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Cournane:2000:BRX

[44] Mavis Cournane. Book review: XML Specification Guide, by Ian S. Graham and Liam Quin. Markup Languages: Theory & Practice, 2(1):107– 108, Winter 2000. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Butler:2000:CTT

[45] Terry Butler, Sue Fisher, Greg Coulombe, Patricia Clements, Isobel Grundy, Susan Brown, Jean Wood, and Rebecca Cameron. Can a team tag consistently? experiences on the Orlando Project. Markup Languages: Theory & Practice, 2(2):111–125, Spring 2000. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

tronic). URL http://mitpress.mit.
edu/catalog/item/default.asp?sid=
3108FA5E-91BB-480F-9930-C1C27725EAB8&
ttype=6&tid=6853.

Howland:2000:REC

[46] Eric Howland and David Niergarth. Regular expressions for checking dates. Markup Languages: Theory & Practice, 2(2):126–132, Spring 2000. CO-DEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic). URL http:

//mitpress.mit.edu/catalog/item/
default.asp?sid=3108FA5E-91BB-480F-
9930-C1C27725EAB8&ttype=6&tid=6855.

Koyanagi:2000:DIX

[47] Teruo Koyanagi, Kouichi Ono, and Masahiro Hori. Demonstrational interface for XSLTM stylesheet generation. Markup Languages: Theory & Practice, 2(2):133–152, Spring 2000. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

tronic). URL http://mitpress.mit.
edu/catalog/item/default.asp?sid=
3108FA5E-91BB-480F-9930-C1C27725EAB8&
ttype=6&tid=6856.

Goldman:2000:SDX

[48] Roy Goldman, Jason McHugh, and Jennifer Widom. From semistructured data to XML. Markup Languages: Theory & Practice, 2(2):153– 163, Spring 2000. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

tronic). URL http://mitpress.mit.
edu/catalog/item/default.asp?sid=
3108FA5E-91BB-480F-9930-C1C27725EAB8&
ttype=6&tid=6857.

Lukka:2000:MPU

[49] Tuomas J. Lukka. Marked-up programming using XML to structure computer program source code. Markup Languages: Theory & Practice, 2(2):165– 182, Spring 2000. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

tronic). URL http://mitpress.mit.
edu/catalog/item/default.asp?sid=
3108FA5E-91BB-480F-9930-C1C27725EAB8&
ttype=6&tid=6858.

Wadler:2000:FSP

[50] Philip Wadler. A formal semantics of patterns in XSLT and XPath. Markup Languages: Theory & Practice, 2(2):183–202, Spring 2000. CO-DEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic). URL http:

//mitpress.mit.edu/catalog/item/
default.asp?sid=3108FA5E-91BB-480F-
9930-C1C27725EAB8&ttype=6&tid=6854.

Smith:2000:MXD

[51] David A. Smith, Anne Mahoney, and Jeffrey A. Rydberg-Cox. Managing XML documents in an integrated digital library. Markup Languages: Theory & Practice, 2(3):205–214, Summer 2000. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

tronic). URL http://mitpress.mit.
edu/catalog/item/default.asp?sid=
3108FA5E-91BB-480F-9930-C1C27725EAB8&
ttype=6&tid=7567.

Sperberg-McQueen:2000:MIM

[52] C. M. Sperberg-McQueen, Claus Huitfeldt, and Allen Renear. Meaning and interpretation of markup. Markup Languages: Theory & Practice, 2(3):215– 234, Summer 2000. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

tronic). URL http://mitpress.mit.
edu/catalog/item/default.asp?sid=
3108FA5E-91BB-480F-9930-C1C27725EAB8&
ttype=6&tid=7568.

Priestley:2000:MWR

[53] Michael Priestley. Managing Web relationships with document structures. Markup Languages: Theory & Practice, 2(3):235–254, Summer 2000. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

tronic). URL http://mitpress.mit.
edu/catalog/item/default.asp?sid=
3108FA5E-91BB-480F-9930-C1C27725EAB8&
ttype=6&tid=7569.

Adler:2000:XMA

[54] Andy Adler, James MacLean, and Alan Boate. An XML messaging architecture for border management systems. Markup Languages: Theory & Practice, 2(3):255–268, Summer 2000. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

tronic). URL http://mitpress.mit.
edu/catalog/item/default.asp?sid=
3108FA5E-91BB-480F-9930-C1C27725EAB8&
ttype=6&tid=7570.

Folch:2000:NTM

[55] Helka Folch, Benoit Habert, and Saadi Lahlou. Navigable Topic Maps for overlaying multiple acquired semantic classifications. Markup Languages: Theory & Practice, 2(3):269–280, Summer 2000. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

tronic). URL http://mitpress.mit.
edu/catalog/item/default.asp?sid=
3108FA5E-91BB-480F-9930-C1C27725EAB8&
ttype=6&tid=7571.

Vorthmann:2000:BS

[56] Scott Vorthmann and Jonathan Robie. Beyond schemas. Markup Languages: Theory & Practice, 2(3):281– 294, Summer 2000. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

tronic). URL http://mitpress.mit.
edu/catalog/item/default.asp?sid=
3108FA5E-91BB-480F-9930-C1C27725EAB8&
ttype=6&tid=7572.

Kimber:2000:UUD

[57] W. Eliot Kimber and John Heintz. Using UML to define XML document types. Markup Languages: Theory & Practice, 2(3):295–320, Summer 2000. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

tronic). URL http://mitpress.mit.
edu/catalog/item/default.asp?sid=
3108FA5E-91BB-480F-9930-C1C27725EAB8&
ttype=6&tid=7573.

Boone:2000:UJX

[58] Keith W. Boone. Using Java for XML processing. Markup Languages: Theory & Practice, 2(3):321–330, Summer 2000. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

tronic). URL http://mitpress.mit.
edu/catalog/item/default.asp?sid=
3108FA5E-91BB-480F-9930-C1C27725EAB8&
ttype=6&tid=7574.

Montour:2000:RDD

[59] Normand Montour. Review: Doc-Book — The Definitive Guide by Norman Walsh and Leonard Muellner. Markup Languages: Theory & Practice, 2(3):330–331, Summer 2000. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

tronic). URL http://mitpress.mit.
edu/catalog/item/default.asp?sid=
3108FA5E-91BB-480F-9930-C1C27725EAB8&
ttype=6&tid=7575.

Bullard:2000:BBG

[60] Len Bullard. Building a better Golem. Markup Languages: Theory & Practice, 2(4):337–351, Fall 2000. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

                                        tronic). URL http://mitpress.mit.
                                        edu/catalog/item/default.asp?sid=
                                        3108FA5E-91BB-480F-9930-C1C27725EAB8&
                                        ttype=6&tid=7576.

Newcomb:2000:TPR

[61] Steven R. Newcomb. Two propositions regarding “Topic Maps”. Markup Languages: Theory & Practice, 2(4): 352, Fall 2000. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

                                        tronic). URL http://mitpress.mit.
                                        edu/catalog/item/default.asp?sid=
                                        3108FA5E-91BB-480F-9930-C1C27725EAB8&
                                        ttype=6&tid=7577.

Schmidt:2000:PNT

[62] Ingrid Schmidt and Carolin Müller. Planning a new type of literary edition: The Thomas Mann Project. Markup Languages: Theory & Practice, 2(4): 353–365, Fall 2000. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

                                        tronic). URL http://mitpress.mit.
                                        edu/catalog/item/default.asp?sid=
                                        3108FA5E-91BB-480F-9930-C1C27725EAB8&
                                        ttype=6&tid=7578.

Rath:2000:TMS

[63] Hans Holger Rath. Topic maps selfcontrol. Markup Languages: Theory & Practice, 2(4):367–388, Fall 2000. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

                                        tronic). URL http://mitpress.mit.
                                        edu/catalog/item/default.asp?sid=
                                        3108FA5E-91BB-480F-9930-C1C27725EAB8&
                                        ttype=6&tid=7579.

Vitali:2000:HFX

[64] Fabio Vitali, Luca Bompani, and Paolo Ciancarini. Hypertext functionalities with XML. Markup Languages: Theory & Practice, 2(4):389– 410, Fall 2000. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

tronic). URL http://mitpress.mit.
edu/catalog/item/default.asp?sid=
3108FA5E-91BB-480F-9930-C1C27725EAB8&
ttype=6&tid=7580.

Renear:2000:DPD

[65] Allen Renear. The descriptive/ procedural distinction is flawed. Markup Languages: Theory & Practice, 2(4): 411–420, Fall 2000. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

tronic). URL http://mitpress.mit.
edu/catalog/item/default.asp?sid=
3108FA5E-91BB-480F-9930-C1C27725EAB8&
ttype=6&tid=7581.

vanWijk:2000:HMF

[66] Diederik A. Gerth van Wijk. How to maintain a family of DTDs and keep them related using switchboard. Markup Languages: Theory & Practice, 2(4): 421–457, Fall 2000. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

tronic). URL http://mitpress.mit.
edu/catalog/item/default.asp?sid=
3108FA5E-91BB-480F-9930-C1C27725EAB8&
ttype=6&tid=7582.

Slotnik:2000:CTU

[67] Arnold M. Slotnik. The Consultant’s Toolkit, updated. Markup Languages: Theory & Practice, 2(4):458, Fall 2000. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

tronic). URL http://mitpress.mit.
edu/catalog/item/default.asp?sid=
3108FA5E-91BB-480F-9930-C1C27725EAB8&
ttype=6&tid=7583.

Tennison:2000:MFX

[68] Jeni Tennison. The many faces of XSLT. Markup Languages: Theory & Practice, 2(4):459–468, Fall 2000. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

                                        tronic). URL http://mitpress.mit.
                                        edu/catalog/item/default.asp?sid=
                                        3108FA5E-91BB-480F-9930-C1C27725EAB8&
                                        ttype=6&tid=7584.

Caton:2001:MCI

[69] Paul Caton. Markup’s current imbalance. Markup Languages: Theory & Practice, 3(1):1–13, Winter 2001. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

                                        tronic). URL http://mitpress.mit.
                                        edu/catalog/item/default.asp?sid=
                                        3108FA5E-91BB-480F-9930-C1C27725EAB8&
                                        ttype=6&tid=7989.

Cover:2001:MTO

[70] Robin Cover. More than one DTD. Markup Languages: Theory & Practice, 3(1):14–16, Winter 2001. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

                                        tronic). URL http://mitpress.mit.
                                        edu/catalog/item/default.asp?sid=
                                        3108FA5E-91BB-480F-9930-C1C27725EAB8&
                                        ttype=6&tid=8002.

Birnbaum:2001:RBG

[71] David J. Birnbaum. The relationship between general and specific DTDs: criticizing TEI critical editions. Markup Languages: Theory & Practice, 3(1):17– 53, Winter 2001. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

                                        tronic). URL http://mitpress.mit.
                                        edu/catalog/item/default.asp?sid=

3108FA5E-91BB-480F-9930-C1C27725EAB8&
ttype=6&tid=7990.

Slotnik:2001:SNGa

[72] Arnold M. Slotnik. SGML: The next generation. Markup Languages: Theory & Practice, 3(1):54, Winter 2001. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

tronic). URL http://mitpress.mit.
edu/catalog/item/default.asp?sid=
3108FA5E-91BB-480F-9930-C1C27725EAB8&
ttype=6&tid=8433.

Herlitz:2001:DXE

[73] Jan Christian Herlitz. The death of XML editors — and the nextgeneration client. Markup Languages: Theory & Practice, 3(1):55–63, Winter 2001. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

tronic). URL http://mitpress.mit.
edu/catalog/item/default.asp?sid=
3108FA5E-91BB-480F-9930-C1C27725EAB8&
ttype=6&tid=7991.

Slotnik:2001:SNGb

[74] Arnold M. Slotnik. SGML: The next generation. Markup Languages: Theory & Practice, 3(1):64, Winter 2001. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

tronic). URL http://mitpress.mit.
edu/catalog/item/default.asp?sid=
3108FA5E-91BB-480F-9930-C1C27725EAB8&
ttype=6&tid=8434.

Marston:2001:OXX

[75] David Marston. OASIS XSLT/Xpath conformance testing. Markup Languages: Theory & Practice, 3(1):65– 71, Winter 2001. CODEN MLTPFG.

ISSN 1099-6621 (print), 1537-2626 (elec-

tronic). URL http://mitpress.mit.
edu/catalog/item/default.asp?sid=
3108FA5E-91BB-480F-9930-C1C27725EAB8&
ttype=6&tid=7992.

Slotnik:2001:SHP

[76] Arnold M. Slotnik. SGML: An historical perspective. Markup Languages: Theory & Practice, 3(1):72, Winter 2001. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

tronic). URL http://mitpress.mit.
edu/catalog/item/default.asp?sid=
3108FA5E-91BB-480F-9930-C1C27725EAB8&
ttype=6&tid=8435.

Hunting:2001:SPS

[77] Sam Hunting. A simple property set for contract architectural forms. Markup Languages: Theory & Practice, 3(1):73– 92, Winter 2001. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

tronic). URL http://mitpress.mit.
edu/catalog/item/default.asp?sid=
3108FA5E-91BB-480F-9930-C1C27725EAB8&
ttype=6&tid=7994.

Liu:2001:PPC

[78] Peiya Liu, Amit Chakraborty, and Liang H. Hsu. Path predicate calculus: Towards a logic formalism for multimedia XML query languages. Markup Languages: Theory & Practice, 3(1):93– 106, Winter 2001. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

tronic). URL http://mitpress.mit.
edu/catalog/item/default.asp?sid=
3108FA5E-91BB-480F-9930-C1C27725EAB8&
ttype=6&tid=7995.

Rizzi:2001:CCF

[79] Romeo Rizzi. Complexity of contextfree grammars with exceptions and the inadequacy of grammars as models for XML and SGML. Markup Languages: Theory & Practice, 3(1):107– 116, Winter 2001. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

tronic). URL http://mitpress.mit.
edu/catalog/item/default.asp?sid=
3108FA5E-91BB-480F-9930-C1C27725EAB8&
ttype=6&tid=7997.

Wood:2001:RBO

[80] Lauren Wood. Review: Building Oracle XML Applications by Steve Muench. Markup Languages: Theory & Practice, 3(1):117–118, Winter 2001. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (elec-

tronic). URL http://mitpress.mit.
edu/catalog/item/default.asp?sid=
3108FA5E-91BB-480F-9930-C1C27725EAB8&
ttype=6&tid=8000.

Mason:2001:FTM

[81] James David Mason. Ferrets and Topic Maps: Knowledge engineering for an analytical engine. Markup Languages: Theory & Practice, 3 (2):123–140, April 2001. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic). URL http:/

/mitpress.mit.eduhttp://mitpress.
mit.edu/catalog/item/default.asp?
sid=81FFBAAF-F78B-4799-8942-1A7B831BFC9A&
ttype=6&tid=8007.

Piez:2001:BDV

[82] Wendell Piez. Beyond the “descriptive vs. procedural” distinction. Markup Languages: Theory & Practice, 3 (2):141–172, April 2001. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic). URL http:/

                                        /mitpress.mit.eduhttp://mitpress.
                                        mit.edu/catalog/item/default.asp?
                                        sid=81FFBAAF-F78B-4799-8942-1A7B831BFC9A&
                                        ttype=6&tid=8011.

Casarini:2001:GDE

[83] Paolo Casarini and Luca Padovani. The Gnome DOM engine. Markup Languages: Theory & Practice, 3 (2):173–190, April 2001. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic). URL http:/

                                        /mitpress.mit.eduhttp://mitpress.
                                        mit.edu/catalog/item/default.asp?
                                        sid=81FFBAAF-F78B-4799-8942-1A7B831BFC9A&
                                        ttype=6&tid=8010.

Abe:2001:VAA

[84] Mari Abe and Masahiro Hori. A visual approach to authoring XPath expressions. Markup Languages: Theory & Practice, 3(2):191–212, April 2001. CO-DEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic). URL http:/

                                        /mitpress.mit.eduhttp://mitpress.
                                        mit.edu/catalog/item/default.asp?
                                        sid=81FFBAAF-F78B-4799-8942-1A7B831BFC9A&
                                        ttype=6&tid=8651.

Brueggemann-Klein:2001:VSC

[85] Anne Brueggemann-Klein, Stefan Hermann, and Derick Wood. The visual specification of context. Markup Languages: Theory & Practice, 3(2):213– 238, April 2001. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Best:2001:OSW

[86] Karl F. Best. OASIS standards work. Markup Languages: Theory & Practice, 3(3):241–249, Summer 2001. CO-DEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic). URL http:/

/mitpress.mit.eduhttp://mitpress.
mit.edu/catalog/item/default.asp?
sid=81FFBAAF-F78B-4799-8942-1A7B831BFC9A&
ttype=6&tid=8426.

Onizuka:2001:XXT

[87] Makoto Onizuka. XTL: An XML transformation language and XSLT generator for XTL. Markup Languages: Theory & Practice, 3(3): 251–284, Summer 2001. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic). URL http:/

/mitpress.mit.eduhttp://mitpress.
mit.edu/catalog/item/default.asp?
sid=81FFBAAF-F78B-4799-8942-1A7B831BFC9A&
ttype=6&tid=8427.

Euzenat:2001:XTF

[88] Jérôme Euzenat and Laurent Tardif. XML transformation flow processing. Markup Languages: Theory & Practice, 3(3):285–311, Summer 2001. CO-DEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic). URL http:/

/mitpress.mit.eduhttp://mitpress.
mit.edu/catalog/item/default.asp?
sid=81FFBAAF-F78B-4799-8942-1A7B831BFC9A&
ttype=6&tid=8428.

Lacher:2001:RTM

[89] Martin S. Lacher and Stefan Decker. RDF, Topic Maps, and the Semantic Web. Markup Languages: Theory & Practice, 3(3):313–331, Summer 2001. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

                                        URL http://mitpress.mit.eduhttp:
                                        //mitpress.mit.edu/catalog/item/
                                        default.asp?sid=81FFBAAF-F78B-4799-
                                        8942-1A7B831BFC9A&ttype=6&tid=8429.

Ogievetsky:2001:XTM

[90] Nikita Ogievetsky. XML Topic Maps through RDF glasses. Markup Languages: Theory & Practice, 3(3): 333–364, Summer 2001. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic). URL http:/

                                        /mitpress.mit.eduhttp://mitpress.
                                        mit.edu/catalog/item/default.asp?
                                        sid=81FFBAAF-F78B-4799-8942-1A7B831BFC9A&
                                        ttype=6&tid=8430.

Liu:2001:PLA

[91] Peiya Liu, Amit Chakraborty, and Liang H. Hsu. A predicate logic approach for MPEG-7 XML document queries. Markup Languages: Theory & Practice, 3(3):365–381, Summer 2001. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

                                        URL http://mitpress.mit.eduhttp:
                                        //mitpress.mit.edu/catalog/item/
                                        default.asp?sid=81FFBAAF-F78B-4799-
                                        8942-1A7B831BFC9A&ttype=6&tid=8431.

Foxwell:2001:RPJ

[92] Harry J. Foxwell. Review: Professional Java Data by Danny Ayres et al. Markup Languages: Theory & Practice, 3(3):382–383, Summer 2001. CO-DEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

Savourel:2001:LX

[93] Yves Savourel. Localization in XML. Markup Languages: Theory & Practice, 3(4):387–393, Fall 2001. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic). URL http:/

/mitpress.mit.eduhttp://mitpress.
mit.edu/catalog/item/default.asp?
sid=81FFBAAF-F78B-4799-8942-1A7B831BFC9A&
ttype=6&tid=9093.

Lubell:2001:AXW

[94] Joshua Lubell. Architectures in an XML world. Markup Languages: Theory & Practice, 3(4):399–410, Fall 2001. CO-DEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic). URL http:/

/mitpress.mit.eduhttp://mitpress.
mit.edu/catalog/item/default.asp?
sid=81FFBAAF-F78B-4799-8942-1A7B831BFC9A&
ttype=6&tid=9094.

Robie:2001:SWS

[95] Jonathan Robie, Lars Marius Garshol, Steve Newcomb, Matthew Fuchs, Libby Miller, Dan Brickley, Vassilis Christophides, and Gregorius Karvounarakis. The Syntactic Web: Syntax and semantics on the Web. Markup Languages: Theory & Practice, 3(4):411–440, Fall 2001. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic). URL http:/

/mitpress.mit.eduhttp://mitpress.
mit.edu/catalog/item/default.asp?
sid=81FFBAAF-F78B-4799-8942-1A7B831BFC9A&
ttype=6&tid=9095.

Piez:2001:XCN

[96] Wendell Piez. The XSL Companion by Neil Bradley. Markup Languages: Theory & Practice, 3 (4):441–445, Fall 2001. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic). URL http:/

/mitpress.mit.eduhttp://mitpress.
mit.edu/catalog/item/default.asp?

                                        sid=81FFBAAF-F78B-4799-8942-1A7B831BFC9A&
                                        ttype=6&tid=9096.

Foxwell:2001:JXE

[97] Harry J. Foxwell. Java and XSLT by Eric M. Burke. Markup Languages: Theory & Practice, 3(4):446–447, Fall 2001. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

                                        URL http://mitpress.mit.eduhttp:
                                        //mitpress.mit.edu/catalog/item/
                                        default.asp?sid=81FFBAAF-F78B-4799-
                                        8942-1A7B831BFC9A&ttype=6&tid=9097.

Hamilton:2001:PXK

[98] Kate Hamilton. Professional XSL by Kurt Cagle et al. Markup Languages: Theory & Practice, 3 (4):449–454, Fall 2001. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic). URL http:/

                                        /mitpress.mit.eduhttp://mitpress.
                                        mit.edu/catalog/item/default.asp?
                                        sid=81FFBAAF-F78B-4799-8942-1A7B831BFC9A&
                                        ttype=6&tid=9098.

Lapeyre:2001:PTU

[99] Deborah A. Lapeyre. Practical Transformation Using XSLT and XPath (XSL Transformations and the XML Path Language) by Crane Softwrights Ltd. Markup Languages: Theory & Practice, 3(4):455–458, Fall 2001. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic). URL http:/

                                        /mitpress.mit.eduhttp://mitpress.
                                        mit.edu/catalog/item/default.asp?
                                        sid=81FFBAAF-F78B-4799-8942-1A7B831BFC9A&
                                        ttype=6&tid=9099.

Gaylord:2001:XQB

[100] Tonya R. Gaylord. XSLT Quickly by Bob DuCharme. Markup Languages:

Theory & Practice, 3(4):458–467, Fall 2001. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).

URL http://mitpress.mit.eduhttp:
//mitpress.mit.edu/catalog/item/
default.asp?sid=81FFBAAF-F78B-4799-
8942-1A7B831BFC9A&ttype=6&tid=9100.

Rosenberg:2001:XEM

[101] Paul Rosenberg. XSL Essentials by Michael Fitzgerald. Markup Languages: Theory & Practice, 3 (4):467–472, Fall 2001. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic). URL http:/

/mitpress.mit.eduhttp://mitpress.
mit.edu/catalog/item/default.asp?
sid=81FFBAAF-F78B-4799-8942-1A7B831BFC9A&
ttype=6&tid=9101.

Gaylord:2001:XWX

[102] Tonya R. Gaylord. XSLT: Working with XML and HTML by Khun Yee Fung. Markup Languages: Theory & Practice, 3(4):472–477, Fall 2001. CO-DEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic). URL http:/

/mitpress.mit.eduhttp://mitpress.
mit.edu/catalog/item/default.asp?
sid=81FFBAAF-F78B-4799-8942-1A7B831BFC9A&
ttype=6&tid=9102.

Lapeyre:2001:XXG

[103] Deborah A. Lapeyre. XSLT & XPath:

A Guide to XML Transformations by John Robert Gardner and Zarella L. Rendon. Markup Languages: Theory & Practice, 3(4):477–480, Fall 2001. CO-DEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic). URL http:/

/mitpress.mit.eduhttp://mitpress.
mit.edu/catalog/item/default.asp?

                                        sid=81FFBAAF-F78B-4799-8942-1A7B831BFC9A&
                                        ttype=6&tid=9103.

Hamilton:2001:XUG

[104] Kate Hamilton. XSLT: The Ultimate Guide to Transforming Web Data by John Hjelm and Peter Stark. Markup Languages: Theory & Practice, 3(4):480–482, Fall 2001. CODEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic). URL http:/

                                        /mitpress.mit.eduhttp://mitpress.
                                        mit.edu/catalog/item/default.asp?
                                        sid=81FFBAAF-F78B-4799-8942-1A7B831BFC9A&
                                        ttype=6&tid=9104.

Hamilton:2001:IXS

[105] Kate Hamilton. Inside XSLT by Steven Holzner. Markup Languages: Theory & Practice, 3(4):482–487, Fall 2001. CO-DEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic). URL http:/

                                        /mitpress.mit.eduhttp://mitpress.
                                        mit.edu/catalog/item/default.asp?
                                        sid=81FFBAAF-F78B-4799-8942-1A7B831BFC9A&
                                        ttype=6&tid=9105.

Lapeyre:2001:XPR

[106] Deborah A. Lapeyre. XSLT: Programmer’s Reference, 2nd Edition by Michael Kay. Markup Languages: Theory & Practice, 3(4):487–492, Fall 2001. CO-DEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic). URL http:/

                                        /mitpress.mit.eduhttp://mitpress.
                                        mit.edu/catalog/item/default.asp?
                                        sid=81FFBAAF-F78B-4799-8942-1A7B831BFC9A&
                                        ttype=6&tid=9106.

Rosenberg:2001:JXJ

[107] Paul Rosenberg. Just XSL by John E. Simpson. Markup Languages: Theory & Practice, 3(4):492–497, Fall 2001. CO-DEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic). URL http:/

/mitpress.mit.eduhttp://mitpress.
mit.edu/catalog/item/default.asp?
sid=81FFBAAF-F78B-4799-8942-1A7B831BFC9A&
ttype=6&tid=9107.

Piez:2001:XXE

[108] Wendell Piez. XSLT and XPath On the Edge, Unlimited Edition by Jeni Tennison. Markup Languages: Theory & Practice, 3(4):498–508, Fall 2001. CO-DEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic). URL http:/

/mitpress.mit.eduhttp://mitpress.
mit.edu/catalog/item/default.asp?
sid=81FFBAAF-F78B-4799-8942-1A7B831BFC9A&
ttype=6&tid=9108.

Piez:2001:XDT

[109] Wendell Piez. XSLT by Doug Tidwell. Markup Languages: Theory & Practice, 3(4):508–513, Fall 2001. CO-DEN MLTPFG. ISSN 1099-6621 (print), 1537-2626 (electronic).